# Critical Filaments and Superconductivity in Quasiperiodic Twisted Bilayer Graphene

### v1

@@ Abstract
Here $\mu^{i}$ is the Pauli matrices acting on the particle-hole space, and $h(\mathbf{r})$ is the qBM Hamiltonian in Eq. (1). The pairing amplitude $\Delta(\mathbf{r})$ is obtained self-consistently by decoupling a local attractive interaction and is spatially inhomogeneous due to quasiperiodicity [30].
Fig. 5 shows the average $\Delta$ and its distribution function for the qBM-BCS model at half-filling obtained with KPM-based mean-field theory [30]. Figs. 5(a) and (b) show the average order parameter $\Delta$ along the two cuts $A_{1}=0.4$ and $1$, respectively. Without the $A_{2}$ potential, the interaction strength is below the critical one for 2D Dirac fermions [42, 43, 44, 45] and $\Delta$ vanishes unless $A_{1}$ is very close to the magic angles. The superconductivity in the qBM model is not fine-tuned as in the BM model, and becomes nonzero as long as $A_{2}$ is not too weak. The strength of the superconductivity correlates with the change of the DOS at the Fermi energy, peaking near filaments with maximal DOS. However, the change in $\Delta$ is much shallower than that of the DOS. It results from the competition of the Fermi energy DOS and the spatial overlap of wave functions. In BCS theory, superconductivity benefits from a larger DOS at the Fermi energy. However, in the BCS-qBM model near the DOS peaks (the filaments), the spatial overlap of wave functions decays faster with energy difference than that in the lakes between filaments (not shown), which makes the wave-function enhancement of superconductivity less significant. Figs. 5(c) and (d) show the distribution of the local order parameter along the two cuts. We can see that the distribution peaks at nonzero values and vanishes at $\Delta_{i}=0$, indicating the superconductivity is nonzero almost everywhere.
@@
Conclusion.—We have shown that the admixture of topology and quasiperiodicity in moiré materials can lead to remarkable quantum critical phenomena and robust superconductivity. Quasiperiodicity acts like disorder for strong mixing, but the coupling of same-valley Dirac fermions for small twists simulates topological surface states. The resulting wave function criticality can strongly boost superconducting pairing without magic-angle fine tuning and without degrading the superfluid stiffness.
While this result applies to twisted bilayer graphene in the chiral limit, we expect distinct multifractality could be found in trilayer systems as well as twisted bilayers with a superimposed, incommensurate superlattice. Even in symmetry classes where disorder localizes in two-dimensions, quasiperiodicity can drive transitions to critical states [48]; this opens up the possibility for similar physics to occur in other van der Waals heterostructures such as transition-metal dichalcogenides where the single-particle physics is simpler.
@@ Acknowledgements.
We thank S. Gopalakrishnan for helpful discussions. This work was supported by the Welch Foundation Grant No. C-1809 (X. Z. and M. S. F.) and by the National Science Foundation under Grant No. DMR-2238895 (J. H. W.). This work was performed in part at Aspen Center for Physics, which is supported by NSF Grant No. PHY-2210452. This work was supported in part by the Big-Data Private-Cloud Research Cyberinfrastructure MRI-award funded by NSF under grant CNS-1338099 and by Rice University’s Center for Research Computing (CRC). Portions of this research were conducted with high performance computing resources provided by Louisiana State University (http://www.hpc.lsu.edu).

### v2

@@ Abstract
The Pauli matrices $\mu^{i}$ act on particle-hole space, and $h(\mathbf{r})$ is the qBM Hamiltonian in Eq. (1). The pairing amplitude $\Delta(\mathbf{r})$ is inhomogeneous due to quasiperiodicity and obtained self-consistently by decoupling a local attractive interaction with coupling strength $U$ [23].
Figure 5: The average superconductivity order parameter $\Delta$, normal-state zero-energy DOS $\nu(0)$, distribution of the local order parameter $\Delta_{i}$ and superfluid stiffness. (a): $\Delta$ and $\nu(E=0)$ along the cut $A_{1}=0.4$; (b): same as (a) along the cut $A_{1}=1$; (c): the distribution of local order parameter $\Delta_{i}$ along $A_{1}=0.4$; (d) same as (c) along the cut $A_{1}=1$; (e): superfluid stiffness (red) along $A_{1}=0.4$ (clean limit value $2\Delta/\pi$ indicated by black dashed curve); (f): same as (e) along $A_{1}=1$. Here $U=4v_{F}k_{\theta}$.
Fig. 5 shows the average $\Delta$ and its distribution function at half-filling obtained with KPM-based mean-field theory [23]. Figs. 5(a) and (b) plot the average order parameter $\Delta$ along the two cuts $A_{1}=0.4$ and $1$, respectively. Without the $A_{2}$ potential, the interaction strength is below the critical one for 2D Dirac fermions [39, 40, 41, 42] and $\Delta$ vanishes unless $A_{1}$ is close to the magic angles. Unlike the BM model, superconductivity in the qBM model is not fine-tuned and becomes nonzero as long as $A_{2}$ is not too weak. The strength of the superconductivity correlates with the change of the DOS at the Fermi energy, peaking near filaments with maximal DOS. However, the change in $\Delta$ is much shallower than that of the DOS. It results from the competition of the Fermi energy DOS and the spatial overlap of normal-state wave functions. In BCS theory, superconductivity benefits from a larger DOS at the Fermi energy. Here however, near the DOS peaks (the filaments), the spatial overlap of wave functions decays faster with energy difference than that in the lakes between filaments (not shown), which reduces the enhancement of superconductivity. Figs. 5(c) and (d) show the distribution of $\Delta_{i}$ and indicate that superconductivity is nonzero almost everywhere.
We compare $\Delta$ in the qBM and usual chiral BM models in Fig. 1. Fig. 1(b) shows $\Delta$ as a function of $U$ in the qBM model for a particular $A_{1,2}$ choice. The scaling indicates BCS-like behavior $\Delta\sim\exp(-c/U)$, where $c$ is a constant, in the $U\rightarrow 0$ limit. This should be contrasted with the chiral BM model, where $\Delta=0$ identically for sufficiently small $U$ away from the magic angle. In [23], we show that finite-energy SQHT states (present due to SWQC) and critical DOS scaling for class CI implies BCS scaling.
@@
Conclusion.—We have shown that the admixture of topology and quasiperiodicity in moiré materials can lead to remarkable quantum critical phenomena and robust superconductivity. Quasiperiodicity acts like disorder for strong mixing, but the coupling of same-valley Dirac fermions for small twists simulates topological surface states. Superconductivity is boosted without magic-angle fine tuning and without degrading the superfluid stiffness.
Taking this toy model seriously, it is worth pointing out that in the microscopic TBLG lattice, our pairing is formally Fulde–Ferrell–Larkin–Ovchinnikov, which would imply a pair-density wave on the lattice scale; further, it would be realized along with the pairing in the other valley to preserve time-reversal. However, the chiral, quasiperiodic BM model we use in Eq. (1) can be exactly constructed in a single layer of graphene with two mismatched $\sqrt{3}$-potentials, as was shown for one such potential in Ref. [45]; in this system, pairing occurs between valleys (and is therefore, s-wave).
The qBM model is a toy version of TTLG, where we expect distinct multifractality will be found [33], as well as in twisted bilayers with a superimposed, incommensurate superlattice. Topology plays an important role; while quasiperiodicity can drive transitions to critical states [46], it may not enhance and could even destroy superconductivity [47]. Nonetheless, this opens up the possibility for similar physics to occur in other moiré materials such as transition-metal dichalcogenide multilayers.
@@ Acknowledgements.
We thank S. Gopalakrishnan and E. König for helpful discussions. We further thank Aaron Dunbrack for pointing out the physical system that realizes our model. This work was supported by the Welch Foundation Grant No. C-1809 (X. Z. and M. S. F.) and by the National Science Foundation under Grant No. DMR-2238895 (J. H. W.). This work was performed in part at Aspen Center for Physics, which is supported by NSF Grant No. PHY-2210452. This work was supported in part by the Big-Data Private-Cloud Research Cyberinfrastructure MRI-award funded by NSF under grant CNS-1338099 and by Rice University’s Center for Research Computing (CRC). Portions of this research were conducted with high performance computing resources provided by Louisiana State University (http://www.hpc.lsu.edu).